# Enhancing Reasoning for Diffusion LLMs via Distribution Matching Policy Optimization

### v1

@@ 5 Conclusion
This paper proposed Distribution Matching Policy Optimization (DMPO), a novel RL fine-tuning framework for dLLMs. DMPO leverages the unique characteristics of dLLMs through importance sampling and WDCE loss, enabling off-policy training and forward-only computation that naturally exploits dLLM inference capabilities. The main limitation of this work is that we focus on a single pretrained dLLM and four elementary reasoning benchmarks, and DMPO’s performance on other pretrained dLLMs and tasks in different domains remains unknown. Our work opens several promising directions for future research, such as investigating the distribution matching framework for other sequence models and studying the design of more effective weight baseline techniques.
@@ Appendix A Related Work
Discrete Diffusion Models.
@@
Diffusion models have been top-performing approaches for generating various data modalities (Zhu et al., 2025f; Esser et al., 2024; Zhu et al., 2025b; Rojas et al., 2025b; Zheng et al., 2025e; Chen et al., 2025a; Ren et al., 2025b). Discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Lou et al., 2024; Zhang et al., 2025b), a natural extension of diffusion models to finite state spaces, have emerged as powerful approaches for generating categorical, sequence data, with applications to text (Nie et al., 2025a; b; Ye et al., 2025), images (Chang et al., 2022; Bai et al., 2025; Shi et al., 2025), and biological sequences (Tang et al., 2025a; Chen et al., 2025b). One of the most effective variants of discrete diffusion models is masked diffusion models (MDM) (Sahoo et al., 2024; Ou et al., 2025; Shi et al., 2024) and its variants (Arriola et al., 2025; Sahoo et al., 2025; Chao et al., 2025). Recently, continuous latents have also been introduced into the modeling of discrete data (Zhang et al., 2025a; Zhou et al., 2025; Zheng et al., 2025c), resulting in improved and more appealing performance.
@@
One particularly important line of development for discrete diffusion models centers on their inference techniques, with the aim of improving generation quality (Nisonoff et al., 2025; Rojas et al., 2025a; Kim et al., 2025; Besnier et al., 2025) and accelerating sampling speed (Ren et al., 2025a; Ben-Hamu et al., 2025; Wu et al., 2025; Hong et al., 2025). Besides these training-free approaches, learning-based approaches, such as few-step distillation, have also achieved decent success for discrete diffusion models (Deschenaux & Gulcehre, 2025; Karimi Monsefi et al., 2025; Zheng et al., 2025b; Zhu et al., 2025d; e). DMPO is closely tied to the literature on fast inference, as it can benefit from it by enjoying a similar training speed acceleration due to its forward nature.
Fine-tuning general discrete diffusion models.
@@
Earlier works on fine-tuning discrete diffusion models primarily focus on applications in biological and chemical domains, e.g., SVDD (Li et al., 2024), DDPP (Rector-Brooks et al., 2025), DRAKES (Wang et al., 2025a), SEPO (Zekri & Boullé, 2025), and TR2-D2 (Tang et al., 2025b). Although these methods work well for their respective tasks, they are not directly applicable to dLLMs due to the unique challenges posed by the language domain, such as large model size, high dimensionality, and the need to maintain linguistic coherence and diversity.
Fine-tuning diffusion LLMs.
@@
Recently, numerous works have proposed RL algorithms for fine-tuning dLLMs, with most existing works being adaptations of the GRPO algorithm (Shao et al., 2024) for AR LLMs. For example, Zhao et al. (2025a) proposed Diffu-GRPO that estimates the per-token response log probabilities via masking all except the required response positions, and partially masking the prompt to get the model output, while their sequence log probability is estimated by mean-field approximation. Gong et al. (2025) introduced Coupled GRPO that modified the Diffu-GRPO method by not partially masking the prompt, and using complementary pairs of masks to mask the same response that fully uses the model output, which we also adopt in our experiments. Yang et al. (2025) proposed UniGRPO, which involves a structured noise strategy and a modified log-likelihood approximation (both per-token and sequence). Concurrent with our work, TraceRL (Wang et al., 2025b) improves dLLM RL training by minimizing a training-inference gap. wd1 (Tang et al., 2025c) introduces additional regularization to the old policy, alongside the regularization applied to the reference model policy, which resembles the case discussed in Sec.˜B.1. We highlight that all these methods are GRPO-style algorithms that require the estimation of per-token response log probabilities, which are typically intractable and challenging for dLLMs. In contrast, our method enjoys the benefit of being a forward one, with more efficiency and accuracy.
Fine-tuning LLMs.
@@
GRPO-style algorithms for fine-tuning diffusion and flow-based models.
@@ Appendix B Theory of Distribution Matching Policy Optimization
B.1 Generalizing WDCE to Zero Temperature with Proximal Descent
@@
Recall that our target distribution is ˜5, which is under a temperature $\alpha>0$. We propose to generalize the WDCE loss ˜11 to incorporate the limiting case $\alpha\to 0$ from the viewpoint of proximal descent (Guo et al., 2025b).
@@
The reward maximization problem ˜4 provides a variational characterization of the target distribution $p_{*}({\bm{o}}|{\bm{q}})$. Suppose now we have a dLLM policy ${\bm{\pi}}_{\theta_{\mathrm{old}}}({\bm{o}}|{\bm{q}})$ that outputs a distribution $p_{\theta_{\mathrm{old}}}({\bm{o}}|{\bm{q}})$. We define the next target distribution $p_{\mathrm{tar}}({\bm{o}}|{\bm{q}})$ as
@@
In fact, the term inside the brackets in ˜17 is $\frac{1}{\eta}\operatorname{KL}(p_{\theta}(\cdot|{\bm{q}})\|p_{\mathrm{tar}}(\cdot|{\bm{q}}))+\operatorname{const}$. This means the next target distribution is a geometric interpolation between the current model distribution $p_{\theta_{\mathrm{old}}}$ and the optimal distribution $p_{*}$, with $\eta>0$ being a step size parameter. ˜18 is well-defined even when $\alpha=0$, although in this case, the target distribution concentrates on the set of maximizers of $r({\bm{q}},{\bm{o}})$ (e.g., all correct question-response pairs) without regularization from the base model $p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})$.
@@
We remark that when picking $\alpha=0$, through the proximal gradient descent formulation, DMPO becomes completely forward-only, as it gets rid of the need for estimating sequence log probability ratio of the form $\log\frac{p_{\text{ref}}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}$, making it the best to incorporate fast dLLM inference techniques for RL training speed-up. However, in this case, we can no longer guarantee the diversity in the target optimal distribution, and thus, we save this direction for future investigation.
B.2 Insights for Weight Baselines: Approximate Variance Reduction

### v2

@@ 5 Conclusion
We propose Distribution Matching Policy Optimization (DMPO), a novel RL fine-tuning framework for dLLMs that leverages the unique characteristics of dLLMs via importance sampling and a WDCE loss, enabling off-policy training and forward-only computation that naturally exploit dLLM inference capabilities. The main limitation of this work is that we focus on a single pretrained dLLM and four elementary reasoning benchmarks, and DMPO’s performance on other pretrained dLLMs and tasks in different domains remains unknown. Our work opens several directions for future research, such as investigating the distribution matching framework for other generative models and studying the design of more effective weight baseline techniques.
@@ Appendix A Related Work
Discrete diffusion models
@@
Diffusion models have been top-performing approaches for generating various data modalities (Zhu et al., 2025f; Esser et al., 2024; Zhu et al., 2025b; Rojas et al., 2025b; Zheng et al., 2025e; Chen et al., 2025a; Ren et al., 2025c). Discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Lou et al., 2024; Zhang et al., 2025b), a natural extension of diffusion models to finite state spaces, have emerged as powerful approaches for generating categorical, sequence data, with applications to text (Nie et al., 2025a, b; Ye et al., 2025), images (Chang et al., 2022; Bai et al., 2025; Shi et al., 2025), and biological sequences (Tang et al., 2025a; Chen et al., 2025b). One of the most effective variants of discrete diffusion models is masked diffusion models (MDM) (Sahoo et al., 2024; Ou et al., 2025; Shi et al., 2024) and its variants (Arriola et al., 2025; Sahoo et al., 2025; Chao et al., 2025). Recently, continuous latents have also been introduced into the modeling of discrete data (Zhang et al., 2025a; Zhou et al., 2025; Zheng et al., 2025c), resulting in improved and more appealing performance.
@@
One particularly important line of development for discrete diffusion models centers on their inference techniques, with the aim of improving generation quality (Nisonoff et al., 2025; Rojas et al., 2025a; Kim et al., 2025; Besnier et al., 2025) and accelerating sampling speed (Ren et al., 2025b; Ben-Hamu et al., 2025; Wu et al., 2025; Hong et al., 2025). Besides these training-free approaches, learning-based approaches, such as few-step distillation, have also achieved decent success for discrete diffusion models (Deschenaux and Gulcehre, 2025; Karimi Monsefi et al., 2025; Zheng et al., 2025b; Zhu et al., 2025d, e). DMPO is closely tied to the literature on fast inference, as it can benefit from it by enjoying a similar training speed acceleration due to its forward nature.
Fine-tuning general discrete diffusion models
@@
Earlier works on fine-tuning discrete diffusion models primarily focus on applications in biological and chemical domains, e.g., SVDD (Li et al., 2025), DDPP (Rector-Brooks et al., 2025), DRAKES (Wang et al., 2025a), SEPO (Zekri and Boullé, 2025), and TR2-D2 (Tang et al., 2025b). Although these methods work well for their respective tasks, they are not directly applicable to dLLMs due to the unique challenges posed by the language domain, such as large model size, high dimensionality, and the need to maintain linguistic coherence and diversity.
Fine-tuning diffusion LLMs
@@
Recently, numerous works have proposed RL algorithms for fine-tuning dLLMs, with most existing works being adaptations of the GRPO algorithm (Shao et al., 2024) for AR LLMs. For example, Zhao et al. (2025a) proposed Diffu-GRPO that estimates the per-token response log probabilities via masking all except the required response positions, and partially masking the prompt to get the model output, while their sequence log probability is estimated by mean-field approximation. Gong et al. (2025) introduced Coupled GRPO that modified the Diffu-GRPO method by not partially masking the prompt, and using complementary pairs of masks to mask the same response that fully uses the model output, which we also adopt in our experiments. Yang et al. (2025) proposed UniGRPO, which involves a structured noise strategy and a modified log-likelihood approximation (both per-token and sequence). Concurrent with our work, TraceRL (Wang et al., 2025b) improves dLLM RL training by minimizing a training-inference gap. wd1 (Tang et al., 2025c) introduces additional regularization to the old policy, alongside the regularization applied to the reference model policy, which resembles the case discussed in Sec. B.2. We highlight that all these methods are GRPO-style algorithms that require estimating per-token response log probabilities, which are typically intractable and challenging for dLLMs. In contrast, our method offers the advantage of being a forward one, with greater efficiency and accuracy.
Fine-tuning LLMs
@@
GRPO-style algorithms for fine-tuning diffusion and flow-based models
@@ Appendix B Theory of Distribution Matching Policy Optimization
B.1 Distribution Matching Policy Optimization from the Stochastic Optimal Control Perspective
This section aims at providing an alternative derivation of DMPO from the perspective of stochastic optimal control (SOC), which is inspired by DRAKES (Wang et al., 2025a) and MDNS (Zhu et al., 2025g). We will first introduce the necessary background on continuous-time Markov chains (CTMCs), then show how MDM sampling can be viewed as a CTMC. Finally, we derive the DMPO framework from the SOC perspective.
Introduction to Continuous-time Markov Chains
To derive the SOC framework for fine-tuning, we view the sampling of an MDM as a time-indexed stochastic process, and the proper mathematical tool is the continuous-time Markov chain (CTMC). A CTMC $X=(X_{t})_{t\in[0,1]}$ is a stochastic process taking value in a discrete state space $\mathcal{X}$. Its law is characterized by the rate matrix $Q=(Q_{t})_{t\in[0,1]}$, defined as
$Q_{t}(x,y)=\lim_{\Delta t\to 0}\frac{\Pr(X_{t+\Delta t}=y|X_{t}=x)-1_{x=y}}{\Delta t},~\forall x,y\in\mathcal{X}.$ (17)
By definition, the off-diagonal entries of $Q_{t}$ are non-negative, and each row sums to zero.
The path of $X$, i.e., $t\mapsto X_{t}(\omega)$, is piecewise constant with discontinuous jumps, and one typically assumes that the path is right continuous with left limits. The path measure a CTMC $X$ is a probability measure on the space of paths defined as $\mathbb{P}^{X}(A):=\Pr(X\in A)$, which is the distribution of $X$. The following lemma shows how to compute the Radon-Nikodým (RN) derivative between two path measures driven by CTMCs with different rate matrices and initial distributions:
Lemma B.1.
Given two CTMCs with rate matrices $Q^{1},Q^{2}$ and initial distributions $\mu_{1},\mu_{2}$ on $\mathcal{X}$, let $\mathbb{P}^{1},\mathbb{P}^{2}$ be the associated path measures. Then, for any path $\xi=(\xi_{t})_{t\in[0,1]}$,
$\log\frac{\mathrm{d}\mathbb{P}^{1}}{\mathrm{d}\mathbb{P}^{2}}(\xi)=\log\frac{\mathrm{d}\mu_{1}}{\mathrm{d}\mu_{2}}(\xi_{0})+\sum_{t:\xi_{t-}\neq\xi_{t}}\log\frac{Q^{1}_{t}(\xi_{t-},\xi_{t})}{Q^{2}_{t}(\xi_{t-},\xi_{t})}+\int_{0}^{1}(Q^{1}_{t}(\xi_{t},\xi_{t})-Q^{2}_{t}(\xi_{t},\xi_{t}))\mathrm{d}t.$ (18)
For the proof, see Campbell et al. (2024, App. C.1), Ren et al. (2025a, Thm. 3.3), or Zhu et al. (2025g, Lem. 1). An intuitive interpretation of 18 is to view the RN derivative as the limit of density ratios between finite-dimensional joint distributions, and approximate the transition probability by 17.
Masked Diffusion Models as Continuous-Time Markov Chains
We will now delve into the CTMC formulation of sampling from an MDM. To avoid notational clutter, we use superscript to denote the position index, and subscript to denote the time index (e.g., ${\bm{\xi}}_{t}=(\xi^{1}_{t},...,\xi^{D}_{t})$). We present the theory only in the case of unconditional generation with sequence length $D$ for simplicity of notations, but it can be easily easily generalized to the case of conditional generation of $\bm{o}$} given a prompt $\bm{q}$}.
As shown in (Ou et al., 2025), by introducing a noise schedule $\gamma(t)=\frac{1}{t}$,  [^3] The choice of noise schedule is essentially not important for MDM. In fact, $\gamma$ can be any positive function with $\int_{0}^{1}\gamma(t)\mathrm{d}t=\infty$. Here, we follow the convention in most of the literature on MDM and choose this specific $\gamma$ such that the conditional distribution of ${\bm{\xi}}_{t}\in\overline{\mathcal{V}}^{D}$ given ${\bm{\xi}}_{1}\in\mathcal{V}^{D}$ is obtained by independently masking each position in ${\bm{\xi}}_{1}$ with probability 1t1-t. the random order autoregressive sampling of an MDM ${\bm{\pi}}_{\theta}$ can be viewed as a CTMC with the rate matrix $Q^{\theta}=(Q^{\theta}_{t})_{t\in[0,1]}$ such that for ${\bm{x}}\neq{\bm{y}}\in\overline{\mathcal{V}}^{D}$,
$Q^{\theta}_{t}({\bm{x}},{\bm{y}})=\gamma(t){\bm{\pi}}_{\theta}({\bm{x}})_{d,n},~\text{if}~{\bm{x}}^{d}=\mathsf{M}~\text{and}~{\bm{y}}={\bm{x}}^{d\leftarrow n},$
and 0 if otherwise, where ${\bm{x}}^{d\leftarrow n}$ means the sequence obtained by replacing the $d$-th position of $\bm{x}$} by $n$. The diagonal terms of $Q^{\theta}_{t}$ can be computed as
$Q^{\theta}_{t}({\bm{x}},{\bm{x}})$ $=-\sum_{{\bm{y}}\neq{\bm{x}}}Q^{\theta}_{t}({\bm{x}},{\bm{y}})=-\sum_{d:{\bm{x}}^{d}=\mathsf{M}}\sum_{n}Q^{\theta}_{t}({\bm{x}},{\bm{x}}^{d\leftarrow n})$
$=-\sum_{d:{\bm{x}}^{d}=\mathsf{M}}\sum_{n}\gamma(t){\bm{\pi}}_{\theta}({\bm{x}})_{d,n}=-\gamma(t)\cdot|\{d:{\bm{x}}^{d}=\mathsf{M}\}|.$ (19)
Therefore, if $\mathbb{P}^{\theta}$, $\mathbb{P}^{\theta^{\prime}}$ are the path measures of the sampling processes of two MDMs parameterized by $\theta$ and $\theta^{\prime}$, respectively, then by 18, assuming that the jump from ${\bm{\xi}}_{t-}$ to ${\bm{\xi}}_{t}$ is at the $d(t)$-th position, we have
$\log\frac{\mathrm{d}\mathbb{P}^{\theta^{\prime}}}{\mathrm{d}\mathbb{P}^{\theta}}({\bm{\xi}})=\sum_{t:{\bm{\xi}}_{t-}\neq{\bm{\xi}}_{t}}\log\frac{{\bm{\pi}}_{\theta^{\prime}}({\bm{\xi}}_{t_{-}})_{d(t),\xi_{t}^{d(t)}}}{{\bm{\pi}}_{\theta}({\bm{\xi}}_{t_{-}})_{d(t),\xi_{t}^{d(t)}}},~\forall{\bm{\xi}}=({\bm{\xi}}_{t})_{t\in[0,1]},$ (20)
as the first term in 18 is always zero (both initial distributions are the point mass on the fully masked sequence), and the diagonal terms in the third term cancel out due to 19.
Moreover, as proved in Ou et al. (2025), the training of an MDM ${\bm{\pi}}_{\theta}$ given i.i.d. samples from the target distribution $p_{\mathrm{data}}$ can be interpreted as minimizing the KL divergence between the target path measure $\mathbb{P}^{*}$ and the parameterized path measure $\mathbb{P}^{\theta}$, where $\mathbb{P}^{*}$ is defined as the path measure of the CTMC with rate matrix $Q^{*}_{t}({\bm{x}},{\bm{x}}^{d\leftarrow n})=\gamma(t)\Pr_{{\bm{X}}\sim p_{\mathrm{data}}}(X^{d}=n|{\bm{X}}^{\mathrm{UM}}={\bm{x}}^{\mathrm{UM}})1_{x^{d}=\mathsf{M}}$, i.e., with the ground-truth conditional distribution. Moreover, one can derive
$\operatorname{KL}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$ $=\operatorname{\mathbb{E}}_{p_{\mathrm{data}}({\bm{x}})}\operatorname{\mathbb{E}}_{m\sim\operatorname{Unif}\{1,...,D\}}\left[\frac{D}{m}\operatorname{\mathbb{E}}_{\mu_{m}({\widetilde{\bm{x}}}|{\bm{x}})}\sum_{d:{\widetilde{\bm{x}}}^{d}=\mathsf{M}}-\log{\bm{\pi}}_{\theta}({\widetilde{\bm{x}}})_{d,x^{d}}\right]+\operatorname{const},$
where $\operatorname{const}$ does not depend on $\theta$, and $\mu_{m}(\cdot|{\bm{x}})$ means to sample a uniformly random subset of $\{1,...,D\}$ of size $m$ and mask the corresponding positions in $\bm{x}$}. Note that this is exactly the denoising cross-entropy loss $\operatorname{\mathbb{E}}_{p_{\mathrm{data}}({\bm{x}})}\mathcal{L}_{\theta}({\bm{x}})$ as presented in 2. In other words, minimizing the KL divergence between sequence-level probabilities $p_{*}({\bm{\xi}}_{1})$ and $p_{\theta}({\bm{\xi}}_{1})\approx\mathrm{e}^{-\mathcal{L}_{\theta}({\bm{\xi}}_{1})}$ in 11 can be interpreted as precisely minimizing the KL divergence between path-level probabilities $\mathbb{P}^{*}({\bm{\xi}})$ and $\mathbb{P}^{\theta}({\bm{\xi}})$.
Fine-tuning MDMs as a Stochastic Optimal Control Problem on Path Measures
The task of fine-tuning a pretrained MDM can be viewed as a stochastic optimal control (SOC) problem on the space of path measures: given a pretrained MDM $\bm{\pi}_{\mathrm{ref}}$ which generates a distribution $p_{\mathrm{ref}}$, we define its induced reference path measure as $\mathbb{P}^{\mathrm{ref}}$, with rate matrix $Q^{\mathrm{ref}}_{t}({\bm{x}},{\bm{x}}^{d\leftarrow n})=\gamma(t)\bm{\pi}_{\mathrm{ref}}({\bm{x}})_{d,n}1_{x^{d}=\mathsf{M}}$, and has terminal distribution $\mathbb{P}^{\mathrm{ref}}_{1}=p_{\mathrm{ref}}$. We aim at finding a target rate matrix $Q^{*}$ such that the associated target path measure $\mathbb{P}^{*}$ has a terminal distribution $p_{*}$ defined in the following way of tilting by reward:
$p_{*}({\bm{x}})$ $=\frac{1}{Z}p_{\mathrm{ref}}({\bm{x}})\mathrm{e}^{r({\bm{x}})/\alpha},~{\bm{x}}\in\mathcal{V}^{D},\qquad\text{where}~Z=\sum_{{\bm{x}}}p_{\mathrm{ref}}({\bm{x}})\mathrm{e}^{r({\bm{x}})/\alpha}.$
This can be achieved by defining the target path measure $\mathbb{P}^{*}$ as
$\mathbb{P}^{*}({\bm{\xi}})=\mathbb{P}^{\mathrm{ref}}({\bm{\xi}}_{[0,1)}|{\bm{\xi}}_{1})p_{*}({\bm{\xi}}_{1})=\mathbb{P}^{\mathrm{ref}}({\bm{\xi}})\frac{p_{*}({\bm{\xi}}_{1})}{p_{\mathrm{ref}}({\bm{\xi}}_{1})}=\frac{1}{Z}\mathbb{P}^{\mathrm{ref}}({\bm{\xi}})\mathrm{e}^{r({\bm{\xi}}_{1})/\alpha},~\forall{\bm{\xi}}=({\bm{\xi}}_{t})_{t\in[0,1]}.$ (21)
We use a network ${\bm{\pi}}_{\theta}$ to parameterize the new rate matrix, initialized at $\bm{\pi}_{\mathrm{ref}}$. Given a current path measure $\mathbb{P}^{\theta}$ induced by a CTMC with rate matrix $Q^{\theta}_{t}({\bm{x}},{\bm{x}}^{d\leftarrow n})=\gamma(t){\bm{\pi}}_{\theta}({\bm{x}})_{d,n}1_{x^{d}=\mathsf{M}}$, we can first derive the RN derivative between the path measures by 20:
$\log\frac{\mathrm{d}\mathbb{P}^{*}}{\mathrm{d}\mathbb{P}^{\theta}}({\bm{\xi}})$ $=\log\frac{\mathrm{d}\mathbb{P}^{*}}{\mathrm{d}\mathbb{P}^{\mathrm{ref}}}({\bm{\xi}})+\log\frac{\mathrm{d}\mathbb{P}^{0}}{\mathrm{d}\mathbb{P}^{\theta}}({\bm{\xi}})$
$=\frac{r({\bm{\xi}}_{1})}{\alpha}-\log Z+\sum_{t:{\bm{\xi}}_{t-}\neq{\bm{\xi}}_{t}}\log\frac{Q^{0}_{t}}{Q^{\theta}_{t}}({\bm{\xi}}_{t-},{\bm{\xi}}_{t})+\int_{0}^{1}\sum_{y\neq{\bm{\xi}}_{t}}(Q^{\theta}_{t}-Q^{0}_{t})({\bm{\xi}}_{t},y)\mathrm{d}t$
$=\frac{r({\bm{\xi}}_{1})}{\alpha}+\sum_{t:{\bm{\xi}}_{t-}\neq{\bm{\xi}}_{t}}\log\frac{\bm{\pi}_{\mathrm{ref}}({\bm{\xi}}_{t-})_{d(t),\xi_{t}^{d(t)}}}{{\bm{\pi}}_{\theta}({\bm{\xi}}_{t-})_{d(t),\xi_{t}^{d(t)}}}-\log Z=:W^{\theta}({\bm{\xi}})-\log Z,$ (22)
where we assume that the jump from ${\bm{\xi}}_{t-}$ to ${\bm{\xi}}_{t}$ is at the $d(t)$-th position. The idea of the weighted denoising cross-entropy (WDCE) loss is essentially to treat i.i.d. samples from the current policy $\mathbb{P}^{\theta}$ as weighted samples from $\mathbb{P}^{*}$, and minimizing the following loss:
$\operatorname{KL}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})+\operatorname{const}$ $=\operatorname{\mathbb{E}}_{p_{*}({\bm{x}})}\mathcal{L}_{\theta}({\bm{x}})=\operatorname{\mathbb{E}}_{\mathbb{P}^{*}({\bm{\xi}})}\mathcal{L}_{\theta}({\bm{\xi}}_{1})$
$=\operatorname{\mathbb{E}}_{\mathbb{P}^{v}({\bm{\xi}})}\frac{\mathrm{d}\mathbb{P}^{*}}{\mathrm{d}\mathbb{P}^{v}}({\bm{\xi}})\mathcal{L}_{\theta}({\bm{\xi}}_{1})=\operatorname{\mathbb{E}}_{\mathbb{P}^{v}({\bm{\xi}})}\frac{1}{Z}\mathrm{e}^{W^{v}({\bm{\xi}})}\mathcal{L}_{\theta}({\bm{\xi}}_{1}),$
where $\mathbb{P}^{v}$ is the path measure induced by a CTMC with rate matrix $Q^{v}$ where the network is parameterized by $v$ (e.g., the old parameters $\theta_{\mathrm{old}}$), whose parameters do not involve gradient calculation. For instance, we can set $v=\theta_{\mathrm{old}}$. Note that $Z=\operatorname{\mathbb{E}}_{\mathbb{P}^{v}({\bm{\xi}})}\mathrm{e}^{W^{v}({\bm{\xi}})}$, which, if estimated via samples, is equivalent to doing softmax normalization on the logits $W^{v}({\bm{\xi}})$ in the batch. Comparing with the WDCE loss 11 presented in Sec. 3.2, we conclude that they are essentially the same.
B.2 Generalizing WDCE to Zero Temperature with Proximal Descent
@@
Recall that our target distribution is 5, which is under a temperature $\alpha>0$. We propose to generalize the WDCE loss 11 to incorporate the limiting case $\alpha\to 0$ from the viewpoint of proximal descent (Guo et al., 2025b).
@@
The reward maximization problem 4 provides a variational characterization of the target distribution $p_{*}({\bm{o}}|{\bm{q}})$. Suppose now we have a dLLM policy ${\bm{\pi}}_{\theta_{\mathrm{old}}}({\bm{o}}|{\bm{q}})$ that outputs a distribution $p_{\theta_{\mathrm{old}}}({\bm{o}}|{\bm{q}})$. We define the next target distribution $p_{\mathrm{tar}}({\bm{o}}|{\bm{q}})$ as
@@
In fact, the term inside the brackets in 23 is $\frac{1}{\eta}\operatorname{KL}(p_{\theta}(\cdot|{\bm{q}})\|p_{\mathrm{tar}}(\cdot|{\bm{q}}))+\operatorname{const}$. This means the next target distribution is a geometric interpolation between the current model distribution $p_{\theta_{\mathrm{old}}}$ and the optimal distribution $p_{*}$, with $\eta>0$ being a step size parameter. 24 is well-defined even when $\alpha=0$, although in this case, the target distribution concentrates on the set of maximizers of $r({\bm{q}},{\bm{o}})$ (e.g., all correct question-response pairs) without regularization from the base model $p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})$.
@@
We remark that when picking $\alpha=0$, through the proximal gradient descent formulation, DMPO becomes completely forward-only, as it eliminates the need for estimating the sequence log probability ratio of the form $\log\frac{p_{\text{ref}}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}$, making it the best option to incorporate fast dLLM inference techniques for RL training speed-up. However, in this case, we can no longer guarantee the diversity in the target optimal distribution, and thus, we save this direction for future investigation.
B.3 Insights for Weight Baselines: Approximate Variance Reduction